OpenStudy (candy13106):

16+4d=8d-4 solve the equation

OpenStudy (mwhit2101):

d=5.

OpenStudy (candy13106):

can u explain

OpenStudy (candy13106):

plz

OpenStudy (mwhit2101):

yes

OpenStudy (candy13106):

ok

OpenStudy (arabpride):

16 + 4d = 8d - 4 -4d -4d 16 = 4d - 4 +4 +4 20 = 4d /4 /4 5 = d or d = 5

OpenStudy (candy13106):

thanks